Suppose that the cost of producing n pairs of shoes is given by the expression C(n)=30+5n. The average cost of producing a pair of shoes is given by A(n)=C(n)/n. What is the amount n for which the average cost 8?

10, 35, 5, 8, 7
Answer #1

Average cost is given as 8

Putting it in the formula of average and total cost , we get

8= C(n)/n => 8n=C(n)

Putting this relation in the expression of cost function , we get

C(n)= 30+5n

8n= 30+5n

3n=30

=> n=10

So number of shoes produced is 10.

(You can comment for doubt)